Imagine that you are viewing a star, located about 100 parsecs away with an apparent magnitude of 0.2. The most likely absolute magnitue of the star is:
a. -4.8
b. 0.1
c. 0.2
d. 0.3

I know the answer is "a" which is -4.8, but I dont know how they got that answer?
1
Expert's answer
2011-02-17T06:27:20-0500
Absolute magnitude can be computed by the formula:
M = m - 5(log10 DL - 1) = 0.2 - 5(log10 100 - 1) = -4.8.